OpenStudy (anonymous):

I need to turn this in to rectangular equation r=4sin(3ø)

OpenStudy (anonymous):

Multiply both side of your equation by r

OpenStudy (anonymous):

then you have to recall and use these relationships: r^2=x^2+y^2 and y=rsin(theta)

OpenStudy (anonymous):

How will multiplying by r help?

OpenStudy (anonymous):

you should get r^2=4rsin(3theta)

OpenStudy (anonymous):

Yes, but I am looking to have x and y and no r and theta